# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $0222\scriptstyle\pm 0.0192 0.0$$067\scriptstyle\pm 0.0058 0.0$$103\scriptstyle\pm 0.0089 0.0$$52\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + KG $0062\scriptstyle\pm 0.0068 0.0$$\scriptstyle\pm 0.0050 0.0055±0.00$$5\scriptstyle\pm 0.0058$ $28\scriptstyle\pm 0.0029 0.0000±0$$\scriptstyle\pm 0.0000 0.000$$000\scriptstyle\pm 0.0000$
L3-FT(Med) $0165\scriptstyle\pm 0.0106 0.0$$083\scriptstyle\pm 0.0058 0.0$$085\scriptstyle\pm 0.0010 0.0$$43\scriptstyle\pm 0.0005 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0094\scriptstyle\pm 0.0082 0.0$$033\scriptstyle\pm 0.0029 0.0$$049\scriptstyle\pm 0.0043 0.0$$25\scriptstyle\pm 0.0021 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(QA) $\mathbf{0.5208}\scriptstyle\pm 0.0525$ $0.4767\scriptstyle\pm 0.0810$ $\mathbf{0.4961}\scriptstyle\pm 0.0677$ $\mathbf{0.3320}\scriptstyle\pm 0.0618$ $0.9667\scriptstyle\pm 0.0577$ $0.9167\scriptstyle\pm 0.0577$
L3-FT(QA) + KG $0.4656\scriptstyle\pm 0.0253$ $0.4217\scriptstyle\pm 0.0584$ $0.4420\scriptstyle\pm 0.0427$ $0.2844\scriptstyle\pm 0.0355$ $0.9333\scriptstyle\pm 0.1155$ $\mathbf{0.9333}\scriptstyle\pm 0.0577$
G-Retriever $0.4563\scriptstyle\pm 0.0579$ $0.3417\scriptstyle\pm 0.0318$ $0.3902\scriptstyle\pm 0.0418$ $0.2433\scriptstyle\pm 0.0313$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
GALAX $0.5046\scriptstyle\pm 0.0833$ $\mathbf{0.4783}\scriptstyle\pm 0.0808$ $0.4909\scriptstyle\pm 0.0819$ $0.3285\scriptstyle\pm 0.0726$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8833\scriptstyle\pm 0.0289$
Table 21: Model performances on STAD
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0057\scriptstyle\pm 0.0100 0.0$$017\scriptstyle\pm 0.0029 0.0$$026\scriptstyle\pm 0.0045 0.0$$13\scriptstyle\pm 0.0023 0.000$$00\scriptstyle\pm 0.0000 0.0$$0167\scriptstyle\pm 0.0289$
L3-FT(Med) + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5528\scriptstyle\pm 0.0311$ $\mathbf{0.5200}\scriptstyle\pm\mathbf{0.0409}$ $\mathbf{0.5353}\scriptstyle\pm\mathbf{0.0325}$ $\mathbf{0.3684}\scriptstyle\pm\mathbf{0.0309}$ $0.9333\scriptstyle\pm 0.0577$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0289}$
L3-FT(QA) + KG $0.5305\scriptstyle\pm 0.0231$ $0.5000\scriptstyle\pm 0.0427$ $0.5144\scriptstyle\pm 0.0332$ $0.3486\scriptstyle\pm 0.0322$ $0.9000\scriptstyle\pm 0.0000$ $0.8667\scriptstyle\pm 0.0289$
G-Retriever $\mathbf{0.5574}\scriptstyle\pm\mathbf{0.0091}$ $0.4883\scriptstyle\pm 0.0029$ $0.5193\scriptstyle\pm 0.0024$ $0.3550\scriptstyle\pm 0.0043$ $0.9000\scriptstyle\pm 0.0000$ $0.9000\scriptstyle\pm 0.0000$
GALAX $0.5298\scriptstyle\pm 0.0331$ $0.4983\scriptstyle\pm 0.0404$ $0.5134\scriptstyle\pm 0.0371$ $0.3465\scriptstyle\pm 0.0338$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$ $0.8667\scriptstyle\pm 0.0289$
Table 22: Model performances on MB
@@
L3 $0333\scriptstyle\pm 0.0577 0.0$$067\scriptstyle\pm 0.0115 0.0$$111\scriptstyle\pm 0.0192 0.0$$57\scriptstyle\pm 0.0098 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + KG $0126\scriptstyle\pm 0.0218 0.0$$067\scriptstyle\pm 0.0115 0.0$$087\scriptstyle\pm 0.0151 0.0$$44\scriptstyle\pm 0.0076 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) $0177\scriptstyle\pm 0.0215 0.0$$100\scriptstyle\pm 0.0100 0.0$$102\scriptstyle\pm 0.0089 0.0$$52\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0282\scriptstyle\pm 0.0237 0.0$$167\scriptstyle\pm 0.0058 0.0$$192\scriptstyle\pm 0.0090 0.0$$097\scriptstyle\pm 0.0046 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
[…]
L3-FT(QA) + KG $0.5086\scriptstyle\pm 0.0633$ $0.5000\scriptstyle\pm 0.0608$ $0.5043\scriptstyle\pm 0.0620$ $0.3387\scriptstyle\pm 0.0569$ $0.7333\scriptstyle\pm 0.2309$ $0.7000\scriptstyle\pm 0.2000$
G-Retriever $\mathbf{0.5546}\scriptstyle\pm\mathbf{0.0421}$ $0.3600\scriptstyle\pm 0.0173$ $0.4365\scriptstyle\pm 0.0257$ $0.2795\scriptstyle\pm 0.0213$ $0.8000\scriptstyle\pm 0.0000$ $0.7000\scriptstyle\pm 0.0000$
GALAX $0.4897\scriptstyle\pm 0.0304$ $0.4867\scriptstyle\pm 0.0351$ $0.4882\scriptstyle\pm 0.0328$ $0.3233\scriptstyle\pm 0.0286$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.8333}\scriptstyle\pm\mathbf{0.0577}$
Table 23: Model performances on ALL
@@
L3 $33\scriptstyle\pm 0.0577 0.0033±0$$3\scriptstyle\pm 0.0058 0.0$$061\scriptstyle\pm 0.0105 0.0$$031\scriptstyle\pm 0.0053 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0015\scriptstyle\pm 0.0026 0.0$$033\scriptstyle\pm 0.0058 0.0$$021\scriptstyle\pm 0.0036 0.0$$10\scriptstyle\pm 0.0018 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0607\scriptstyle\pm 0.0528 0.0$$133\scriptstyle\pm 0.0115 0.0$$219\scriptstyle\pm 0.0189 0.0$$111\scriptstyle\pm 0.0096 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5487\scriptstyle\pm 0.0561$ $0.5400\scriptstyle\pm 0.0625$ $0.5443\scriptstyle\pm 0.0592$ $0.3754\scriptstyle\pm 0.0569$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.1155}$
L3-FT(QA) + KG $0.6010\scriptstyle\pm 0.0608$ $0.5600\scriptstyle\pm 0.1100$ $0.5786\scriptstyle\pm 0.0868$ $0.4106\scriptstyle\pm 0.0872$ $0.9333\scriptstyle\pm 0.1155$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.1155}$
G-Retriever $0.6285\scriptstyle\pm 0.0043$ $0.5300\scriptstyle\pm 0.0000$ $0.5750\scriptstyle\pm 0.0018$ $0.4036\scriptstyle\pm 0.0018$ $0.8000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.6860}\scriptstyle\pm\mathbf{0.0246}$ $\mathbf{0.6700}\scriptstyle\pm\mathbf{0.0265}$ $\mathbf{0.6779}\scriptstyle\pm\mathbf{0.0255}$ $\mathbf{0.5131}\scriptstyle\pm\mathbf{0.0295}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0577}$
Table 24: Model performances on LGG

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $0222\scriptstyle\pm 0.0192 0.0$$067\scriptstyle\pm 0.0058 0.0$$103\scriptstyle\pm 0.0089 0.0$$52\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $0062\scriptstyle\pm 0.0068 0.0$$\scriptstyle\pm 0.0050 0.0055±0.00$$5\scriptstyle\pm 0.0058$ $28\scriptstyle\pm 0.0029 0.0000±0$$\scriptstyle\pm 0.0000 0.000$$000\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0165\scriptstyle\pm 0.0106 0.0$$083\scriptstyle\pm 0.0058 0.0$$085\scriptstyle\pm 0.0010 0.0$$43\scriptstyle\pm 0.0005 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0094\scriptstyle\pm 0.0082 0.0$$033\scriptstyle\pm 0.0029 0.0$$049\scriptstyle\pm 0.0043 0.0$$25\scriptstyle\pm 0.0021 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5208\scriptstyle\pm 0.0525$ $0.4767\scriptstyle\pm 0.0810$ $0.4961\scriptstyle\pm 0.0677$ $0.3320\scriptstyle\pm 0.0618$ $0.9667\scriptstyle\pm 0.0577$ $0.9167\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics + KG $0.4656\scriptstyle\pm 0.0253$ $0.4217\scriptstyle\pm 0.0584$ $0.4420\scriptstyle\pm 0.0427$ $0.2844\scriptstyle\pm 0.0355$ $0.9333\scriptstyle\pm 0.1155$ $0.9333\scriptstyle\pm 0.0577$
G-Retriever + pre-GAT $0.4563\scriptstyle\pm 0.0579$ $0.3417\scriptstyle\pm 0.0318$ $0.3902\scriptstyle\pm 0.0418$ $0.2433\scriptstyle\pm 0.0313$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
RoG $0.5038\scriptstyle\pm 0.0245$ $0.4550\scriptstyle\pm 0.0954$ $0.4717\scriptstyle\pm 0.0730$ $0.3123\scriptstyle\pm 0.0583$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9667\scriptstyle\pm 0.0289$
SubgraphRAG $\mathbf{0.5360}\scriptstyle\pm\mathbf{0.0133}$ $0.4550\scriptstyle\pm 0.0433$ $0.4869\scriptstyle\pm 0.0223$ $0.3237\scriptstyle\pm 0.0186$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
GNN-RAG $0.4854\scriptstyle\pm 0.0257$ $0.4717\scriptstyle\pm 0.0375$ $0.4782\scriptstyle\pm 0.0319$ $0.3164\scriptstyle\pm 0.0263$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8833\scriptstyle\pm 0.0577$
GALAX $0.5046\scriptstyle\pm 0.0833$ $0.4783\scriptstyle\pm 0.0808$ $0.4909\scriptstyle\pm 0.0819$ $0.3285\scriptstyle\pm 0.0726$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8833\scriptstyle\pm 0.0289$
GALAX (Qwen2.5-7B) $0.5237\scriptstyle\pm 0.0212$ $\mathbf{0.5150}\scriptstyle\pm\mathbf{0.0180}$ $\mathbf{0.5193}\scriptstyle\pm\mathbf{0.0196}$ $\mathbf{0.3509}\scriptstyle\pm\mathbf{0.0178}$ $0.9333\scriptstyle\pm 0.0577$ $0.8500\scriptstyle\pm 0.0500$
Table 29: Model performance on STAD
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0057\scriptstyle\pm 0.0100 0.0$$017\scriptstyle\pm 0.0029 0.0$$026\scriptstyle\pm 0.0045 0.0$$13\scriptstyle\pm 0.0023 0.000$$00\scriptstyle\pm 0.0000 0.0$$0167\scriptstyle\pm 0.0289$
L3-FT(Med) + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5528\scriptstyle\pm 0.0311$ $0.5200\scriptstyle\pm 0.0409$ $0.5353\scriptstyle\pm 0.0325$ $0.3684\scriptstyle\pm 0.0309$ $0.9333\scriptstyle\pm 0.0577$ $0.9333\scriptstyle\pm 0.0289$
L3-FT(CRISPR) + Omics + KG $0.5305\scriptstyle\pm 0.0231$ $0.5000\scriptstyle\pm 0.0427$ $0.5144\scriptstyle\pm 0.0332$ $0.3486\scriptstyle\pm 0.0322$ $0.9000\scriptstyle\pm 0.0000$ $0.8667\scriptstyle\pm 0.0289$
G-Retriever + pre-GAT $0.5574\scriptstyle\pm 0.0091$ $0.4883\scriptstyle\pm 0.0029$ $0.5193\scriptstyle\pm 0.0024$ $0.3550\scriptstyle\pm 0.0043$ $0.9000\scriptstyle\pm 0.0000$ $0.9000\scriptstyle\pm 0.0000$
RoG $0.5099\scriptstyle\pm 0.0498$ $0.4950\scriptstyle\pm 0.0606$ $0.5020\scriptstyle\pm 0.0556$ $0.3392\scriptstyle\pm 0.0475$ $0.8000\scriptstyle\pm 0.2646$ $0.7167\scriptstyle\pm 0.2021$
SubgraphRAG $0.5044\scriptstyle\pm 0.0443$ $0.4950\scriptstyle\pm 0.0606$ $0.4993\scriptstyle\pm 0.0531$ $0.3353\scriptstyle\pm 0.0437$ $67\scriptstyle\pm 0.2309 0.7667±0$$67\scriptstyle\pm 0.2309$
GNN-RAG $0.5387\scriptstyle\pm 0.0068$ $0.5267\scriptstyle\pm 0.0029$ $0.5325\scriptstyle\pm 0.0047$ $0.3680\scriptstyle\pm 0.0037$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8167\scriptstyle\pm 0.0577$
GALAX $0.5298\scriptstyle\pm 0.0331$ $0.4983\scriptstyle\pm 0.0404$ $0.5134\scriptstyle\pm 0.0371$ $0.3465\scriptstyle\pm 0.0338$ $0.9667\scriptstyle\pm 0.0577$ $0.8667\scriptstyle\pm 0.0289$
GALAX (Qwen2.5-7B) $0.5731\scriptstyle\pm 0.0132$ $\mathbf{0.5883}\scriptstyle\pm\mathbf{0.0029}$ $\mathbf{0.5803}\scriptstyle\pm\mathbf{0.0055}$ $\mathbf{0.4117}\scriptstyle\pm\mathbf{0.0057}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9833}\scriptstyle\pm\mathbf{0.0289}$
Table 30: Model performance on MB
@@
L3 + Omics $0333\scriptstyle\pm 0.0577 0.0$$067\scriptstyle\pm 0.0115 0.0$$111\scriptstyle\pm 0.0192 0.0$$56\scriptstyle\pm 0.0098 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $0126\scriptstyle\pm 0.0218 0.0$$067\scriptstyle\pm 0.0115 0.0$$087\scriptstyle\pm 0.0151 0.0$$44\scriptstyle\pm 0.0076 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0177\scriptstyle\pm 0.0215 0.0$$100\scriptstyle\pm 0.0100 0.0$$102\scriptstyle\pm 0.0089 0.0$$52\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0282\scriptstyle\pm 0.0237 0.0$$167\scriptstyle\pm 0.0058 0.0$$192\scriptstyle\pm 0.0090 0.0$$097\scriptstyle\pm 0.0046 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics $8\scriptstyle\pm 0.0327 0.5033±0.05$$3\scriptstyle\pm 0.0551$ $0.5164\scriptstyle\pm 0.0440$ $0.3488\scriptstyle\pm 0.0394$ $0.8667\scriptstyle\pm 0.1155$ $0.8000\scriptstyle\pm 0.1000$
L3-FT(CRISPR) + Omics + KG $0.5086\scriptstyle\pm 0.0633$ $0.5000\scriptstyle\pm 0.0608$ $0.5043\scriptstyle\pm 0.0620$ $0.3387\scriptstyle\pm 0.0569$ $0.7333\scriptstyle\pm 0.2309$ $0.7000\scriptstyle\pm 0.2000$
G-Retriever + pre-GAT $\mathbf{0.5546}\scriptstyle\pm\mathbf{0.0421}$ $0.3600\scriptstyle\pm 0.0173$ $0.4365\scriptstyle\pm 0.0257$ $0.2794\scriptstyle\pm 0.0213$ $0.8000\scriptstyle\pm 0.0000$ $0.7000\scriptstyle\pm 0.0000$
RoG $0.5530\scriptstyle\pm 0.0211$ $\mathbf{0.5400}\scriptstyle\pm\mathbf{0.0173}$ $\mathbf{0.5464}\scriptstyle\pm\mathbf{0.0192}$ $\mathbf{0.3761}\scriptstyle\pm\mathbf{0.0183}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9000\scriptstyle\pm 0.0000$
SubgraphRAG $0.5545\scriptstyle\pm 0.0119$ $0.3900\scriptstyle\pm 0.1212$ $0.4502\scriptstyle\pm 0.0737$ $0.2925\scriptstyle\pm 0.0632$ $0.8667\scriptstyle\pm 0.1155$ $0.9000\scriptstyle\pm 0.0000$
GNN-RAG $0.5306\scriptstyle\pm 0.0267$ $0.4467\scriptstyle\pm 0.1097$ $0.4776\scriptstyle\pm 0.0606$ $0.3151\scriptstyle\pm 0.0512$ $0.9333\scriptstyle\pm 0.1155$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0577}$
GALAX $0.4897\scriptstyle\pm 0.0304$ $0.4867\scriptstyle\pm 0.0351$ $0.4882\scriptstyle\pm 0.0328$ $0.3233\scriptstyle\pm 0.0286$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.8333\scriptstyle\pm 0.0577$
GALAX (Qwen2.5-7B) $0.5300\scriptstyle\pm 0.0346$ $0.5300\scriptstyle\pm 0.0346$ $0.5300\scriptstyle\pm 0.0346$ $0.3611\scriptstyle\pm 0.0325$ $0.8667\scriptstyle\pm 0.1155$ $0.9333\scriptstyle\pm 0.0577$
Table 31: Model performance on ALL
@@
L3 + Omics $33\scriptstyle\pm 0.0577 0.0033±0$$3\scriptstyle\pm 0.0058 0.0$$061\scriptstyle\pm 0.0105 0.0$$031\scriptstyle\pm 0.0053 0.0$$667\scriptstyle\pm 0.1155 0.0$$0333\scriptstyle\pm 0.0577$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0015\scriptstyle\pm 0.0026 0.0$$033\scriptstyle\pm 0.0058 0.0$$021\scriptstyle\pm 0.0036 0.0$$10\scriptstyle\pm 0.0018 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0607\scriptstyle\pm 0.0528 0.0$$133\scriptstyle\pm 0.0115 0.0$$219\scriptstyle\pm 0.0189 0.0$$111\scriptstyle\pm 0.0096 0.0$$\scriptstyle\pm 0.0000 0.0000±0.00$$0\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5487\scriptstyle\pm 0.0561$ $0.5400\scriptstyle\pm 0.0624$ $0.5443\scriptstyle\pm 0.0592$ $0.3754\scriptstyle\pm 0.0569$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.1155$
L3-FT(CRISPR) + Omics + KG $0.6010\scriptstyle\pm 0.0608$ $0.5600\scriptstyle\pm 0.1100$ $0.5786\scriptstyle\pm 0.0868$ $0.4106\scriptstyle\pm 0.0872$ $0.9333\scriptstyle\pm 0.1155$ $0.9333\scriptstyle\pm 0.1155$
G-Retriever + pre-GAT $0.6285\scriptstyle\pm 0.0043$ $0.5300\scriptstyle\pm 0.0000$ $0.5750\scriptstyle\pm 0.0018$ $0.4036\scriptstyle\pm 0.0018$ $0.8000\scriptstyle\pm 0.0000$ $0.8000\scriptstyle\pm 0.0000$
RoG $0.5596\scriptstyle\pm 0.0699$ $0.5133\scriptstyle\pm 0.1501$ $0.5326\scriptstyle\pm 0.1168$ $0.3684\scriptstyle\pm 0.1042$ $0.9333\scriptstyle\pm 0.1155$ $0.9667\scriptstyle\pm 0.0577$
SubgraphRAG $0.6000\scriptstyle\pm 0.0000$ $0.6000\scriptstyle\pm 0.0000$ $0.6000\scriptstyle\pm 0.0000$ $0.4286\scriptstyle\pm 0.0000$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
GNN-RAG $0.5682\scriptstyle\pm 0.0858$ $0.5200\scriptstyle\pm 0.1609$ $0.5402\scriptstyle\pm 0.1293$ $0.3770\scriptstyle\pm 0.1186$ $0.9333\scriptstyle\pm 0.1155$ $0.9667\scriptstyle\pm 0.0577$
GALAX $\mathbf{0.6860}\scriptstyle\pm\mathbf{0.0246}$ $\mathbf{0.6700}\scriptstyle\pm\mathbf{0.0265}$ $\mathbf{0.6779}\scriptstyle\pm\mathbf{0.0255}$ $\mathbf{0.5131}\scriptstyle\pm\mathbf{0.0295}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9333\scriptstyle\pm 0.0577$
GALAX (Qwen2.5-7B) $0.6000\scriptstyle\pm 0.0520$ $0.6000\scriptstyle\pm 0.0520$ $0.6000\scriptstyle\pm 0.0520$ $0.4299\scriptstyle\pm 0.0542$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$
Table 32: Model performance on LGG